# Disambiguating Reference in Visually Grounded Dialogues through Joint Modeling of Textual and Multimodal Semantic Structures

### v1

@@ Abstract
Multimodal reference resolution, including phrase grounding, aims to understand the semantic relations between mentions and real-world objects. Phrase grounding between images and their captions is a well-established task. In contrast, for real-world applications, it is essential to integrate textual and multimodal reference resolution to unravel the reference relations within dialogue, especially in handling ambiguities caused by pronouns and ellipses. This paper presents a framework that unifies textual and multimodal reference resolution by mapping mention embeddings to object embeddings and selecting mentions or objects based on their similarity [^1] The code is publicly available at https://github.com/SInadumi/mmrr.. Our experiments show that learning textual reference resolution, such as coreference resolution and predicate-argument structure analysis, positively affects performance in multimodal reference resolution. In particular, our model with coreference resolution performs better in pronoun phrase grounding than representative models for this task, MDETR and GLIP. Our qualitative analysis demonstrates that incorporating textual reference relations strengthens the confidence scores between mentions, including pronouns and predicates, and objects, which can reduce the ambiguities that arise in visually grounded dialogues.
@@
Shun Inadumi 1,2, Nobuhiro Ueda 3,*, Koichiro Yoshino 4,2,1, 1 Nara Institute of Science and Technology, 2 Guardian Robot Project, RIKEN 3 Kyoto University, 4 Institute of Science Tokyo inazumi.shun.in6@naist.ac.jp ueda@nlp.ist.i.kyoto-u.ac.jp yoshino@c.titech.ac.jp
@@ 1 Introduction
Pronouns and ellipses are typical instances of direct and indirect reference. These are frequently used in dialogue and pose challenges for multimodal reference resolution Ueda et al. (2024). Phrase grounding between images and their captions is a well-established task Kazemzadeh et al. (2014); Plummer et al. (2017); however, identifying direct references is still insufficient in visually grounded dialogues [^2] Example of Table 3 shows the performance of GLIP Li et al. (2022) for phrase grounding in Japanese dialogue. The Recall@1 is 0.377, significantly worse than that for general Japanese captions.. Figure 1 illustrates an example where pronouns and ellipses of topical terms, such as subjects and objects, present difficulties for such a framework.
@@ 2.1.2 Textual Reference Resolution
Given a text $\mathbf{T}$, TRR identifies phrases that have a reference relation with another phrase. We refer to such phrases as mentions. We use the term TRR to refer collectively to coreference resolution, predicate-argument structure (PAS) analysis Iida et al. (2007), and bridging anaphora (BA) resolution Poesio and Vieira (1998); Yu and Poesio (2020).
@@ 2.2 J-CRe3 Annotation
Ambiguous expressions, including pronouns and ellipses, frequently occur in spoken language Piantadosi et al. (2012), with ellipses occurring especially in Japanese Seki et al. (2002), Chinese Kong and Zhou (2010), and Korean Park et al. (2015). In the ellipses, for example, indirect references from predicates to objects can exist without explicitly mentioning those objects. Following the previous work Ueda et al. (2024), we refer to these instances as zero references, as in the case of zero anaphora in text Sasano et al. (2008). While datasets addressing pronouns have existed in the past Das et al. (2017); Kottur et al. (2021); Wu et al. (2023); Goel et al. (2023b), J-CRe3 differs in that it also includes zero references [^5] Although the work of Oguz et al. (2023, 2024) addresses noun phrase ellipsis, J-CRe3 is the only dataset that provides annotations for the explicit estimation of zero references from predicates to objects..
@@ 3.2.1 Textual Reference Resolution Model
The embeddings $\mathbf{T}^{\prime}$ are at a subword level, but a mention is at a basic phrase [^6] This phrase consists of one content word and zero or more function words. level. To link these different units, we use the first subword of a mention as its main representation during learning and inference. This step is the same as for our MRR model described below.
@@ 4.1 Settings
For the TRR, each training instance consists of three sentences, shifting by one sentence at a time. For phrase grounding and MRR, the instance unit definitions vary by dataset. In J-CRe3, each training instance comprises three utterances, shifting by one utterance at a time, paired with an image; evaluation is performed on individual utterance–image pairs [^7] See Appendix C for ablation study results on the utterance length of $\mathbf{T}$ during evaluation of phrase grounding and MRR.. In Flickr30k-Ent-JP, each instance, used for training and evaluation, consists of up to five captions paired with an image.

### v2

@@ Abstract
Multimodal reference resolution, including phrase grounding, aims to understand the semantic relations between mentions and real-world objects. Phrase grounding between images and their captions is a well-established task. In contrast, for real-world applications, it is essential to integrate textual and multimodal reference resolution to unravel the reference relations within dialogue, especially in handling ambiguities caused by pronouns and ellipses. This paper presents a framework that unifies textual and multimodal reference resolution by mapping mention embeddings to object embeddings and selecting mentions or objects based on their similarity. [^1] The code is publicly available at https://github.com/SInadumi/mmrr. Our experiments show that learning textual reference resolution, such as coreference resolution and predicate-argument structure analysis, positively affects performance in multimodal reference resolution. In particular, our model with coreference resolution performs better in pronoun phrase grounding than representative models for this task, MDETR and GLIP. Our qualitative analysis demonstrates that incorporating textual reference relations strengthens the confidence scores between mentions, including pronouns and predicates, and objects, which can reduce the ambiguities that arise in visually grounded dialogues.
@@
Shun Inadumi 1,2, Nobuhiro Ueda 3,*, Koichiro Yoshino 4,2,1, 1 Nara Institute of Science and Technology, 2 Guardian Robot Project, RIKEN 3 Kyoto University, 4 Institute of Science Tokyo inazumi.shun.in6@naist.ac.jp ueda@nlp.ist.i.kyoto-u.ac.jp koichiro@c.titech.ac.jp
@@ 1 Introduction
Pronouns and ellipses are typical instances of direct and indirect reference. These are frequently used in dialogue and pose challenges for multimodal reference resolution Ueda et al. (2024). Phrase grounding between images and their captions is a well-established task Kazemzadeh et al. (2014); Plummer et al. (2017); however, identifying direct references is still insufficient in visually grounded dialogues. [^2] Example of Table 3 shows the performance of GLIP Li et al. (2022) for phrase grounding in Japanese dialogue. The Recall@1 is 0.377, significantly worse than that for general Japanese captions. Figure 1 illustrates an example where pronouns and ellipses of topical terms, such as subjects and objects, present difficulties for such a framework.
@@ 2.1.2 Textual Reference Resolution
Given a text $\mathbf{T}$, TRR identifies phrases that have a reference relation with another phrase. We refer to such phrases as mentions. We use the term TRR to refer collectively to coreference resolution, predicate-argument structure (PAS) analysis Iida et al. (2007), and bridging anaphora (BA) resolution Poesio and Vieira (1998); Kobayashi and Ng (2020); Yu and Poesio (2020).
@@ 2.2 J-CRe3 Annotation
Ambiguous expressions, including pronouns and ellipses, frequently occur in spoken language Piantadosi et al. (2012), with ellipses occurring especially in Japanese Seki et al. (2002), Chinese Kong and Zhou (2010), and Korean Park et al. (2015). In the ellipses, for example, indirect references from predicates to objects can exist without explicitly mentioning those objects. Following the previous work Ueda et al. (2024), we refer to these instances as zero references, as in the case of zero anaphora in text Sasano et al. (2008). While datasets addressing pronouns have existed in the past Das et al. (2017); Kottur et al. (2021); Wu et al. (2023); Goel et al. (2023b), J-CRe3 differs in that it also includes zero references. [^5] Although the work of Oguz et al. (2023, 2024) addresses noun phrase ellipsis, J-CRe3 is the only dataset that provides annotations for the explicit estimation of zero references from predicates to objects.
@@ 3.2.1 Textual Reference Resolution Model
The embeddings $\mathbf{T}^{\prime}$ are at a subword level, but a mention is at a basic phrase level. [^6] This phrase consists of one content word and zero or more function words. To link these different units, we use the first subword of a mention as its main representation during learning and inference. This step is the same as for our MRR model described below.
@@ 4.1 Settings
For the TRR, each training instance consists of three sentences, shifting by one sentence at a time. For phrase grounding and MRR, the instance unit definitions vary by dataset. In J-CRe3, each training instance comprises three utterances, shifting by one utterance at a time, paired with an image; evaluation is performed on individual utterance–image pairs. [^7] See Appendix C for ablation study results on the utterance length of $\mathbf{T}$ during evaluation of phrase grounding and MRR. In Flickr30k-Ent-JP, each instance, used for training and evaluation, consists of up to five captions paired with an image.